# Contribution of remote bands to orbital magnetization in twisted bilayer graphene

### 04 Mar 2026 v1

@@ IV Orbital magnetization
where $n$ denotes the band index, $\epsilon_{\mu\nu}$ is the antisymmetric tensor, $\mu,\nu$ are Cartesian indices, and $\partial_{\mu}\equiv\partial/\partial k_{\mu}$. For the orbital magnetization, it has been recently shown Kang et al. (2025) that both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ can be evaluated using the HF Hamiltonian and its eigenfunctions Souza and Vanderbilt (2008); Resta (2020); Ji et al. (2025):
$M_{orb}=\frac{e}{2i\hbar}\sum_{n\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\left|\left(H^{HF}(\bm{k})+\epsilon_{n}^{HF}(\bm{k})-2\mu\right)\right|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}(\epsilon_{n}^{HF}(\bm{k})-\mu)$ (11)
$m_{SR}=\frac{e}{2i\hbar}\sum_{n,\alpha}\sum_{\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\middle|u_{\alpha}(\bm{k})\right\rangle\left(\epsilon^{HF}_{\alpha}(\bm{k})-\epsilon_{n}^{HF}(\bm{k})\right)\left\langle u_{\alpha}(\bm{k})\middle|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}\left(\epsilon_{n}^{HF}(\bm{k})-\mu\right)\left(1-n_{F}(\epsilon_{\alpha}^{HF}(\bm{k})-\mu)\right)$ (12)
@@
We emphasize that the Hartree-Fock Hamiltonian $H^{\mathrm{HF}}(\mathbf{k})$ must be formulated in the full Hilbert space, including the flat-band sector as well as both the occupied and empty remote bands. In principle, Eqs. 11 and  12 involve summations over all occupied states in both the active and remote bands, rendering the calculation formally infinite. To make the numerical calculations tractable, we evaluate $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ in a truncated Hilbert space. For clarity of notation, we label the HF Bloch states $|u_{n}^{A/R+/R-}(\mathbf{k})\rangle$ by the superscripts $A$, $R+$, and $R$-, which denote the active bands, the empty remote bands, and the occupied remote bands, respectively. The empty and occupied remote bands are indexed by $n$, counted outward from the charge neutrality point, such that the states $|u_{n}^{R+}(\mathbf{k})\rangle$ and $|u_{n}^{R-}(-\mathbf{k})\rangle$ are related by the particle-hole symmetry Bernevig et al. (2021). On this basis, we introduce two projection operators onto the truncated occupied and empty states.
@@
Because the projection operators $P_{n_{cut}}(\mathbf{k})$ and $Q_{n_{cut}}(\mathbf{k})$ are manifestly gauge invariant, this formulation avoids the gauge ambiguities that typically arise in direct evaluations of orbital magnetization based on Bloch wave functions. Moreover, it circumvents the singularities associated with band degeneracy points. Formally, physical quantities are defined in the limit $n_{\mathrm{cut}}\to\infty$. Nevertheless, we anticipate rapid convergence with increasing $n_{\mathrm{cut}}$, enabling reliable numerical evaluation within a finite truncated Hilbert space.
IV.1 $\nu=\pm 3$
We first consider the case of fillings $\nu=\pm 3$. The self-consistent Hartree–Fock calculation yields a spin–valley polarized Chern insulating ground state characterized by spontaneous breaking of the combined $C_{2}\mathcal{T}$ symmetry. The corresponding band dispersions, shown in Fig. 2(a) and (b), indicate that among the eight active bands, a single pair associated with one spin–valley flavor develops a gap due to $C_{2}\mathcal{T}$ symmetry breaking and acquires a finite Chern number. The remaining three pairs of bands, corresponding to the other spin–valley flavors (shown in black), remain topological trivial. The symmetry breaking generates an indirect energy gap of approximately 13 meV between the occupied Chern band and the higher-energy bands. These states correspond to the experimentally observed quantum anomalous Hall phases with Chern number $C=1$ (C=1C=-1) at fillings $\nu=3$ ($\nu=-3$Sharpe et al. (2019); Serlin et al. (2020); Nuckolls et al. (2020).
@@ IV.1 $\nu=\pm 3$
We next compute the orbital magnetization of these states using Eqs. 15 and 16. As shown in Fig.2(c,d), both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation component $m_{\mathrm{SR}}$ converge for $n_{\mathrm{cut}}\approx 20$. Moreover, the results respect the particle–hole symmetry relating fillings $\nu=3$ and $\nu=-3$, satisfying $M_{\mathrm{orb}}(\nu=3)=M_{\mathrm{orb}}(\nu=-3)$, and $m_{\mathrm{SR}}(\nu=3)=-m_{\mathrm{SR}}(\nu=-3)$ sm , consistent with the experimentally observed Landau fan structureDas et al. (2021); Nuckolls et al. (2020). Since convergence is achieved for $n_{\mathrm{cut}}\approx 20$, we adopt $n_{\mathrm{cut}}=30$ in the following calculations to ensure numerical stability.
One might be tempted to conclude that both the orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ arise exclusively from the occupied Chern bands that explicitly break time-reversal symmetry, and that only these bands need to be considered—analogous to the case of Berry curvature. Such reasoning, however, is not generally valid. Although the occupied states of the remote bands, $|u_{n}(\mathbf{k})\rangle$, remain invariant under time-reversal symmetry, their momentum derivatives $|\partial_{\mu}u_{n}(\mathbf{k})\rangle$ generally possess non-vanishing overlaps with those states in the active bands. Consequently, once time-reversal symmetry is broken within the active bands, the nominally time-reversal-symmetric remote bands acquire finite contributions to both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$. The need to include a substantial number of remote bands—at least $20$ pairs in our calculations—to achieve numerical convergence provides direct evidence of their non-negligible contributions.
More physically relevant is the behavior of $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ within the Chern insulating gap, indicated by the red dots in Fig. 3(a,b). The total orbital magnetization $M_{\mathrm{orb}}$ exhibits the expected linear dependence on $\mu$. Thermodynamics gives the relation $\partial M_{orb}/\partial\mu=\partial N/\partial B$, consistent with the Diophantine equation that characterizes quantum Hall states. Taking the derivative of Eq.11 with respect to $\mu$ yields
@@
Figure 3: Calculated total orbital magnetization $M_{\mathrm{orb}}$ (a) and self-rotation contribution $m_{\mathrm{SR}}$ (b) per moiré cell as functions of the chemical potential for the Chern insulating state at $\nu=3$ inside the insulating gap, whose band dispersion is shown in Fig. 2(a). $M_{\mathrm{orb}}$ shows a linear dependence with $\mu$ and changes sign within the gap while $m_{\mathrm{SR}}$ remains constant. The number of remote band pairs $n_{cut}$ is set to 30 in the calculation.
IV.2 Competing states at $\nu=\pm 1$
@@
We next consider the correlated insulating states at fillings $\nu=\pm 1$. Owing to the particle–hole symmetry relating $\nu$ and $\nu$, we focus on the case $\nu=1$. The Hartree–Fock calculation yields two nearly degenerate correlated insulating states. The lowest-energy state is a mixed phase combining intervalley coherence (IVC) and $C_{2}\mathcal{T}$ symmetry breaking. In this state, the five valence bands consist of two bands associated with a spin-polarized IVC sector carrying zero Chern number, one $C_{2}\mathcal{T}$-broken band with finite Chern number, and an additional pair of ungapped bands from one spin–valley flavor, resulting in a total Chern number $C=1$ Zhang et al. (2020). Its band dispersion is shown in Fig.4(b). This state is lower in energy by approximately 0.5 meV per particle compared to the state in which three valence bands break $C_{2}\mathcal{T}$ symmetry and each carries a finite Chern number. We refer to the latter as the correlated Chern insulator (CCI) with total Chern number $C=3$, whose dispersion is shown in Fig.4(a). We then compute the orbital magnetization for both states. As shown in Fig.4(c), the total orbital magnetization $M_{\mathrm{orb}}$ for both states exhibits a linear dependence on the chemical potential within the insulating gap and changes sign across the gap. Both the magnitude and the slope of $M_{\mathrm{orb}}$ are larger for the CCI than for the mixed IVC state, reflecting its higher Chern number. This implies that the CCI is energetically favored in the presence of a finite magnetic field, consistent with experimental observations. Similarly, the self-rotation contribution $m_{\mathrm{SR}}$ of the CCI is significantly larger than that of the IVC state, as shown in Fig.4(d).
@@
Figure 4: (a,b) Band dispersions of the correlated Chern insulating state (CCI) (a) and the intervalley-coherent (IVC) state (b) at filling $\nu=1$, obtained within the HF approximation. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). For the CCI, the Chern bands are triply degenerate, yielding a total Chern number $C=3$. (c,d) Calculated total orbital magnetization $M_{\mathrm{orb}}$ (c) and self-rotation contribution $m_{\mathrm{SR}}$ (d) per moiré cell as functions of the chemical potential $\mu$ inside the insulating gap for the two states shown in (a) and (b). Solid circles correspond to the CCI, and solid squares to the IVC state. $M_{\mathrm{orb}}$ varies linearly with $\mu$ and changes sign across the gap, while $m_{\mathrm{SR}}$ remains constant. The number of remote band pairs $n_{cut}$ is set to 30 in the calculation.

### 16 Apr 2026 v2

@@ IV Orbital magnetization
where $n$ denotes the band index, $\epsilon_{\mu\nu}$ is the antisymmetric tensor, $\mu,\nu$ are Cartesian indices, and $\partial_{\mu}\equiv\partial/\partial k_{\mu}$. For the orbital magnetization, it has been recently proved Kang et al. (2025) that both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ can be evaluated using the HF Hamiltonian and its eigenfunctions Souza and Vanderbilt (2008); Resta (2020); Ji et al. (2025):
$M_{orb}=\frac{e}{2i\hbar}\sum_{n\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\left|\left(H^{\rm HF}(\bm{k})+E_{n}^{\rm HF}(\bm{k})-2\mu\right)\right|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}\left(E_{n}^{\rm HF}(\bm{k})-\mu\right)$ (6)
$m_{SR}=\frac{e}{2i\hbar}\sum_{n,\alpha}\sum_{\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\middle|u_{\alpha}(\bm{k})\right\rangle\left(E^{\rm HF}_{\alpha}(\bm{k})-E_{n}^{\rm HF}(\bm{k})\right)\left\langle u_{\alpha}(\bm{k})\middle|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}\left(E_{n}^{\rm HF}(\bm{k})-\mu\right)\left(1-n_{F}\left(E_{\alpha}^{\rm HF}(\bm{k})-\mu\right)\right)$ (7)
@@
For orbital magnetization, it is important to emphasize that the Hartree-Fock Hamiltonian $H^{\mathrm{HF}}(\mathbf{k})$ entering Eqs. 6 and 7 is, strictly speaking, defined in the full Hilbert space, including the active flat-band sector together with both the occupied and empty remote bands. The derivation in Ref. Kang et al. (2025) assumes that the HF problem is solved self-consistently in this full Hilbert space. In the present work, however, the HF calculation is carried out only within the active flat-band subspace, while the remote bands are kept fixed. This frozen-band treatment may therefore introduce quantitative errors into the resulting orbital magnetization. A fully self-consistent $B=0$ HF calculation including additional remote bands, and a careful study of the convergence of the orbital magnetization with respect to that extension, are left for future work.
Even within this approximation, Eqs. 6 and 7 still involve sums over infinitely many occupied and empty remote states. To make the calculation tractable, we evaluate $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ in a truncated Hilbert space. For clarity of notation, we label the HF Bloch states $|u_{n}^{A/R+/R-}(\mathbf{k})\rangle$ by the superscripts $A$, $R+$, and $R$-, which denote the active bands, the empty remote bands, and the occupied remote bands, respectively. The remote bands are indexed outward from charge neutrality, such that the states $|u_{n}^{R+}(\mathbf{k})\rangle$ and $|u_{n}^{R-}(-\mathbf{k})\rangle$ are related by particle-hole symmetry Bernevig et al. (2021). On this basis, we introduce two projection operators onto the truncated occupied and empty sectors:
@@
Because the projection operators $P_{n_{cut}}(\mathbf{k})$ and $Q_{n_{cut}}(\mathbf{k})$ are manifestly gauge invariant, this formulation avoids the gauge ambiguities that typically arise in direct evaluations of orbital magnetization based on $\bm{k}$-derivatives of Bloch wave functions. Moreover, it circumvents the singularities associated with band degeneracy points. Formally, physical quantities are defined in the limit $n_{\mathrm{cut}}\to\infty$. Nevertheless, we anticipate rapid convergence with increasing $n_{\mathrm{cut}}$, enabling reliable numerical evaluation within a finite truncated Hilbert space.
IV.1 Valley magnetization for the BM model
We first apply the projection-matrix formalism to the non-interacting BM model. To generate a finite valley magnetization, we introduce a staggered potential through an additional term $\Delta_{C2B}\sigma_{z}$ in the BM Hamiltonian. This term breaks $C_{2}$ symmetry and thus gaps the Dirac points, as shown in Fig. 2(a). Because the staggered potential preserves time-reversal symmetry, the orbital magnetizations in the two valleys remain equal in magnitude and opposite in sign, such that the total orbital magnetization vanishes. We therefore restrict our analysis to the $\tau=$- valley and evaluate $M_{\mathrm{orb}}$ within this sector. The self-rotation contribution $m_{\mathrm{SR}}$ is not evaluated in this case, as it is logarithmic divergent in the ultraviolet limit unless an additional high-energy cutoff is introduced sm .
To highlight the importance of the truncation procedure, we compare our scheme with an alternative prescription, denoted m1, in which only the occupied projection matrix $P_{n_{\rm cut}}(\mathbf{k})$ is truncated, while the empty-state projection matrix is taken as $Q_{n_{\rm cut}}^{\prime}(\mathbf{k})=I-P_{n_{\rm cut}}(\mathbf{k})$. As shown in Fig. 2(b), truncating both projection matrices yields smooth and rapid convergence of $M_{\mathrm{orb}}$ with increasing $n_{\mathrm{cut}}$ (black curves). In contrast, scheme m1 exhibits substantial oscillations and attains convergence only when essentially all remote bands are included in the occupied projection matrix, where the two prescriptions become equivalent. The origin of this oscillatory behavior and the numerical stability of the symmetric truncation scheme are discussed in the Supplemental Material sm .
Figure 2: (a) Band structure for valley ($\tau=$-) in the non-interacting BM model with a staggered potential $\Delta_{C2B}=20$ meV. The gapped low-energy flat bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). (b) Calculated total orbital magnetization $M_{\mathrm{orb}}$ per moiré cell for the valley - state shown in (a), obtained using two truncation schemes within the projection-matrix formalism and plotted as functions of the number of remote-band pairs $n_{\mathrm{cut}}$. When evaluating $M_{\mathrm{orb}}$, the chemical potential is fixed at $\mu=-10$ meV, inside the gap. The inset shows the enlarged view of panel (b), highlighting the oscillatory behavior of scheme m1.
IV.2 $\nu=\pm 3$
With the truncation scheme validated in the non-interacting BM model, we now apply the same formalism to the interacting Hartree-Fock states, beginning with the Chern insulating phases at $\nu=\pm 3$. The self-consistent Hartree–Fock calculation yields a spin–valley polarized Chern insulating ground state characterized by spontaneous breaking of the combined $C_{2}\mathcal{T}$ symmetry. The corresponding band dispersions, shown in Fig. 3(a) and (b), indicate that among the eight active bands, a single pair associated with one spin–valley flavor develops a gap due to $C_{2}\mathcal{T}$ symmetry breaking and acquires a finite Chern number. The remaining three pairs of bands, corresponding to the other spin–valley flavors (shown in black), remain topological trivial. The symmetry breaking generates an indirect energy gap of approximately 13 meV between the occupied Chern band and the higher-energy bands. These states correspond to the experimentally observed quantum anomalous Hall phases with Chern number $C=1$ (C=1C=-1) at fillings $\nu=3$ ($\nu=-3$Sharpe et al. (2019); Serlin et al. (2020); Nuckolls et al. (2020).
@@ IV.2 $\nu=\pm 3$
We next evaluate the orbital magnetization of these states using Eqs. 10 and 11. As shown in Fig. 3(c,d), both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ converge by $n_{\mathrm{cut}}\approx 20$. The results also respect the particle-hole symmetry relating fillings $\nu=3$ and $\nu=-3$, with $M_{\mathrm{orb}}(\nu=3)=M_{\mathrm{orb}}(\nu=-3)$ and $m_{\mathrm{SR}}(\nu=3)=-m_{\mathrm{SR}}(\nu=-3)$ sm , consistent with the experimentally observed Landau-fan structure Das et al. (2021); Nuckolls et al. (2020). Since convergence is already achieved for $n_{\mathrm{cut}}\approx 20$, we use $n_{\mathrm{cut}}=30$ in the following calculations to ensure numerical stability.
One might be tempted to conclude that both the orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ arise exclusively from the occupied Chern bands that explicitly break time-reversal symmetry, and that only these bands need to be considered, in analogy with the case of Berry curvature. Such reasoning, however, is not generally valid. Although the occupied states of the remote bands, $|u_{n}(\mathbf{k})\rangle$, remain invariant under time-reversal symmetry, their momentum derivatives $|\partial_{\mu}u_{n}(\mathbf{k})\rangle$ generally possess non-vanishing overlaps with those states in the active bands. Consequently, once time-reversal symmetry is broken within the active bands, the nominally time-reversal-symmetric remote bands acquire finite contributions to both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$. The need to include at least $20$ remote-band pairs to reach numerical convergence provides direct evidence that these contributions are substantial.
To make the role of the remote bands more transparent, we fix the HF band structure obtained at $\nu=3$ and introduce an auxiliary energy level $\mu^{\prime}$ in place of the physical chemical potential $\mu$. We then vary $\mu^{\prime}$ rigidly from 40-40 meV, well below the active HF bands, to $80$ meV, well above them, and compute both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ as functions of $\mu^{\prime}$. For a fixed HF spectrum, the resulting $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ represent the contributions from all active bands below $\mu^{\prime}$ together with the negative-energy remote bands. As shown in Fig. 4(a,b), the blue plateaus in both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ for $\mu^{\prime}\lesssim-20$ meV arise entirely from the remote bands and reach magnitudes as large as $\sim 7.5\mu_{B}$, directly demonstrating their sizable contribution to the orbital magnetization. In this regime, $M_{\mathrm{orb}}=m_{\mathrm{SR}}$ because the remote bands are topologically trivial and carry zero Berry curvature.
More physically relevant is the behavior of $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ within the Chern insulating gap, indicated by the red dots in Fig. 4(a,b). Inside the Chern gap, the auxiliary energy level $\mu^{\prime}$ is equivalent to the true chemical potential $\mu$ of the Chern insulating state. The total orbital magnetization $M_{\mathrm{orb}}$ exhibits the expected linear dependence on $\mu$. Thermodynamics gives the relation $\partial M_{orb}/\partial\mu=\partial N/\partial B$, consistent with the Diophantine equation that characterizes quantum Hall states. Taking the derivative of Eq.6 with respect to $\mu$ yields
@@
Figure 4: Calculated total orbital magnetization $M_{\mathrm{orb}}$ (a) and self-rotation contribution $m_{\mathrm{SR}}$ (b) per moiré cell as functions of the energy level $\mu^{\prime}$ for the Chern insulating state at $\nu=3$ inside the insulating gap, whose band dispersion is shown in Fig. 3(a). The red dots denote the values within the Chern insulating gap, where $M_{\mathrm{orb}}$ shows a linear dependence with $\mu^{\prime}$($\mu$) and changes sign within the gap while $m_{\mathrm{SR}}$ remains constant. The blue dots represent the contribution from the remote bands arising from the time-reversal symmetry breaking of the HF Hamiltonian $H_{\mathrm{tot}}^{\mathrm{HF}}$. The number of remote band pairs $n_{cut}$ is set to 30 in the calculation.
IV.3 Competing states at $\nu=\pm 1$
@@
We next consider the correlated insulating states at fillings $\nu=\pm 1$. Owing to the particle–hole symmetry relating $\nu$ and $\nu$, we focus on the case $\nu=1$. The Hartree–Fock calculation yields two types of nearly degenerate correlated insulating states. The lowest-energy state is a mixed phase combining intervalley coherence (IVC) and $C_{2}\mathcal{T}$ symmetry breaking. Its band dispersion is shown in Fig. 5(b). In this state, the five valence bands consist of two bands associated with a spin-polarized IVC sector carrying zero Chern number, one $C_{2}\mathcal{T}$-broken band with finite Chern number, and an additional pair of ungapped bands from a single spin-valley flavor, yielding a total Chern number $C=1$ Zhang et al. (2020). This state is lower in energy by approximately 1.52 meV per moiré cell than a family of states in which three valence bands break $C_{2}\mathcal{T}$ symmetry and each carry a finite Chern number. We refer to the latter as correlated Chern insulators (CCIs). Among these degenerate CCI solutions, we present the state with the largest Chern number, namely $C=3$, whose dispersion is shown in Fig. 5(a).
We then evaluate the orbital magnetization for both states. As shown in Fig. 5(c), the total orbital magnetization $M_{\mathrm{orb}}$ depends linearly on the chemical potential within the insulating gap and changes sign across the gap for both states. Both the magnitude and the slope of $M_{\mathrm{orb}}$ are larger for the CCI than for the mixed IVC state, reflecting its larger Chern number. This implies that the CCI is energetically favored in the presence of a finite magnetic field, consistent with experimental observations. The self-rotation contribution $m_{\mathrm{SR}}$, shown in Fig. 5(d), is likewise significantly larger in magnitude for the CCI than for the mixed IVC state. Using the difference in orbital magnetization at the center of the gap, $\delta M_{\mathrm{orb}}=4.31\,\mu_{B}$ per moiré cell, together with the energy difference between the two insulating states, $\delta E=1.52$ meV per moiré cell, we estimate a critical magnetic field for the transition from the mixed IVC state to the CCI of $B_{c}\sim\delta E/\delta M_{\mathrm{orb}}\approx 6.1$ T. We also note a sign-convention subtlety in comparison with experiment. In Landau-fan analyses based on the Streda formula, the Chern number is extracted from the slope of the carrier density with respect to the applied magnetic field. Because the charge carriers are electrons and therefore carry negative charge, the physical Chern number is the negative of this slope, in agreement with the convention used in our theoretical analysis. Thus, for direct comparison with the experimentally reported values, one should consider the time-reversed partner of the ground state obtained in our calculation, under which both the Chern number and the associated orbital magnetization change sign.
@@
Figure 5: (a,b) HF band dispersions of the correlated Chern insulating state (CCI) (a) and the mixed intervalley-coherent (IVC) state (b) at filling $\nu=1$. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). For the CCI, the Chern bands are triply degenerate, yielding a total Chern number $C=3$. (c,d) Total orbital magnetization $M_{\mathrm{orb}}$ (c) and self-rotation contribution $m_{\mathrm{SR}}$ (d) per moiré cell as functions of the chemical potential $\mu$ inside the insulating gap for the two states shown in (a) and (b). Solid circles correspond to the CCI, and solid squares to the mixed IVC state. $M_{\mathrm{orb}}$ varies linearly with $\mu$ and changes sign across the gap, while $m_{\mathrm{SR}}$ remains constant. The number of remote-band pairs is fixed at $n_{\mathrm{cut}}=30$ in the calculation.